A chemist needs to mix 75% salwater solution with a 50% saltwater solution to obtain 10 gallons of a 60% saltwater solution. How many gallons of each of the solutions must be used?
Answer by rfer(10425)

.75x+.5(10-x)=.6(10)
.75x+5-.5x=6
.25x=1
x=4 gal of 75%
x-4=6 gal of 50%